In a regression analysis, if SST = 4500 and SSE = 1575, then the coefficient of determination is _____.
Answer #1

Solution :

Given that,

SST = 4500

SSE = 1575

then the coefficient of determination is = 1 - (SSE / SST) = 1 - (1575 / 4500) = 0.65